OpenStudy (danielpreece):

A regular hexagon has sides measuring 6 centimeters. The hexagon is enlarged using a scale factor of 3 over 2. What are the lengths of each side of the enlarged hexagon?

jhonyy9 (jhonyy9):

6*3/2 =18/2 =9

OpenStudy (danielpreece):

Thank you!

OpenStudy (lia22):

6*3/2 =18/2 =9

OpenStudy (mathstudent55):

3/2 = 1.5 Each side becomes 1.5 times longer. 1.5 * 6 = 9